OpenStudy (anonymous):

find three consecutive odd integers whose sum is 195... how do u solve it?

OpenStudy (anonymous):

ok, you're solving for three numbers, so you should start by converting them to variables.. ill use x, y z. Following so far?

OpenStudy (anonymous):

yes...continue!

OpenStudy (anonymous):

haha ok, well with these variables we can write this equation x+ y + z = 195 After that, you can convert y and z into terms of x by noticing they are consecutive odd integers. any idea how to do that?

OpenStudy (anonymous):

no...no idea...

OpenStudy (anonymous):

alright, since these numbers are consecutive and odd, y = x+2 and z = y+2 = x+2+2, can you see how that works? so if x was 1, y would be 3, z would be 5

OpenStudy (anonymous):

o...im followin...

OpenStudy (anonymous):

so with the original equation, x + y + z = 195 you substitute the new values to get x+ (x+2) +(x+2+2) = 195 can you solve for x?

OpenStudy (anonymous):

can i guess and check?!?!?!?

OpenStudy (anonymous):

haha that wouldnt work very well, you have to solve 3x + 6 = 195 by first subtracting 6 from both sides then dividing by 3. What do you get?

OpenStudy (anonymous):

63!!!!so x=63,y=65,and z=67!!!and 63+65+67=195!!!! yes!!!thanks a million!!!

OpenStudy (anonymous):

yup you got it :)